\begin{thebibliography}{27}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Anandkumar et~al.(2014)Anandkumar, Ge, Hsu, Kakade, and
  Telgarsky]{anandkumar2014tensor}
Animashree Anandkumar, Rong Ge, Daniel Hsu, Sham~M Kakade, and Matus Telgarsky.
\newblock Tensor decompositions for learning latent variable models.
\newblock \emph{The Journal of Machine Learning Research}, 15\penalty0
  (1):\penalty0 2773--2832, 2014.

\bibitem[Barak and Moitra(2016)]{barak2016noisy}
Boaz Barak and Ankur Moitra.
\newblock Noisy tensor completion via the sum-of-squares hierarchy.
\newblock In \emph{Conference on Learning Theory}, pages 417--445, 2016.

\bibitem[Barak et~al.(2019)Barak, Hopkins, Kelner, Kothari, Moitra, and
  Potechin]{barak2019nearly}
Boaz Barak, Samuel Hopkins, Jonathan Kelner, Pravesh~K Kothari, Ankur Moitra,
  and Aaron Potechin.
\newblock A nearly tight sum-of-squares lower bound for the planted clique
  problem.
\newblock \emph{SIAM Journal on Computing}, 48\penalty0 (2):\penalty0 687--735,
  2019.

\bibitem[Berthet and Rigollet(2013)]{berthet2013complexity}
Quentin Berthet and Philippe Rigollet.
\newblock Complexity theoretic lower bounds for sparse principal component
  detection.
\newblock In \emph{Conference on Learning Theory}, pages 1046--1066, 2013.

\bibitem[Bollob{\'a}s and Erd{\"o}s(1976)]{bollobas1976cliques}
B{\'e}la Bollob{\'a}s and Paul Erd{\"o}s.
\newblock Cliques in random graphs.
\newblock In \emph{Mathematical Proceedings of the Cambridge Philosophical
  Society}, volume~80, pages 419--427. Cambridge University Press, 1976.

\bibitem[Brennan and Bresler(2020)]{brennan2020reducibility}
Matthew Brennan and Guy Bresler.
\newblock Reducibility and statistical-computational gaps from secret leakage.
\newblock \emph{arXiv preprint arXiv:2005.08099}, 2020.

\bibitem[Brennan et~al.(2018)Brennan, Bresler, and
  Huleihel]{brennan2018reducibility}
Matthew Brennan, Guy Bresler, and Wasim Huleihel.
\newblock Reducibility and computational lower bounds for problems with planted
  sparse structure.
\newblock In \emph{Conference On Learning Theory}, pages 48--166, 2018.

\bibitem[Dekel et~al.(2014)Dekel, Gurel-Gurevich, and Peres]{dekel2014finding}
Yael Dekel, Ori Gurel-Gurevich, and Yuval Peres.
\newblock Finding hidden cliques in linear time with high probability.
\newblock \emph{Combinatorics, Probability and Computing}, 23\penalty0
  (1):\penalty0 29--49, 2014.

\bibitem[Deshpande and Montanari(2015{\natexlab{a}})]{deshpande2015finding}
Yash Deshpande and Andrea Montanari.
\newblock Finding hidden cliques of size $\sqrt{N/e}$ in nearly linear time.
\newblock \emph{Foundations of Computational Mathematics}, 15\penalty0
  (4):\penalty0 1069--1128, 2015{\natexlab{a}}.

\bibitem[Deshpande and Montanari(2015{\natexlab{b}})]{deshpande2015improved}
Yash Deshpande and Andrea Montanari.
\newblock Improved sum-of-squares lower bounds for hidden clique and hidden
  submatrix problems.
\newblock In \emph{Conference on Learning Theory}, pages 523--562,
  2015{\natexlab{b}}.

\bibitem[Feldman et~al.(2017)Feldman, Grigorescu, Reyzin, Vempala, and
  Xiao]{feldman2017statistical}
Vitaly Feldman, Elena Grigorescu, Lev Reyzin, Santosh~S Vempala, and Ying Xiao.
\newblock Statistical algorithms and a lower bound for detecting planted
  cliques.
\newblock \emph{Journal of the ACM (JACM)}, 64\penalty0 (2):\penalty0 8, 2017.

\bibitem[Gamarnik and Zadik(2019)]{gamarnik2019landscape}
David Gamarnik and Ilias Zadik.
\newblock The landscape of the planted clique problem: Dense subgraphs and the
  overlap gap property.
\newblock \emph{arXiv preprint arXiv:1904.07174}, 2019.

\bibitem[Hajek et~al.(2015)Hajek, Wu, and Xu]{hajek2015computational}
Bruce Hajek, Yihong Wu, and Jiaming Xu.
\newblock Computational lower bounds for community detection on random graphs.
\newblock In \emph{Conference on Learning Theory}, pages 899--928, 2015.

\bibitem[Hillar and Lim(2013)]{hillar2013most}
Christopher~J Hillar and Lek-Heng Lim.
\newblock Most tensor problems are np-hard.
\newblock \emph{Journal of the ACM (JACM)}, 60\penalty0 (6):\penalty0 45, 2013.

\bibitem[Hopkins(2018)]{hopkins2018statistical}
Samuel Hopkins.
\newblock Statistical inference and the sum of squares method.
\newblock \emph{Cornell thesis}, 2018.

\bibitem[Hopkins et~al.(2015)Hopkins, Shi, and Steurer]{hopkins2015tensor}
Samuel~B Hopkins, Jonathan Shi, and David Steurer.
\newblock Tensor principal component analysis via sum-of-square proofs.
\newblock In \emph{Conference on Learning Theory}, pages 956--1006, 2015.

\bibitem[Jerrum(1992)]{jerrum1992large}
Mark Jerrum.
\newblock Large cliques elude the metropolis process.
\newblock \emph{Random Structures \& Algorithms}, 3\penalty0 (4):\penalty0
  347--359, 1992.

\bibitem[Kearns(1998)]{kearns1998efficient}
Michael Kearns.
\newblock Efficient noise-tolerant learning from statistical queries.
\newblock \emph{Journal of the ACM (JACM)}, 45\penalty0 (6):\penalty0
  983--1006, 1998.

\bibitem[Knowles and Yin(2013)]{knowles2013isotropic}
Antti Knowles and Jun Yin.
\newblock The isotropic semicircle law and deformation of wigner matrices.
\newblock \emph{Communications on Pure and Applied Mathematics}, 66\penalty0
  (11):\penalty0 1663--1749, 2013.

\bibitem[Lesieur et~al.(2017)Lesieur, Miolane, Lelarge, Krzakala, and
  Zdeborov{\'a}]{lesieur2017statistical}
Thibault Lesieur, L{\'e}o Miolane, Marc Lelarge, Florent Krzakala, and Lenka
  Zdeborov{\'a}.
\newblock Statistical and computational phase transitions in spiked tensor
  estimation.
\newblock In \emph{2017 IEEE International Symposium on Information Theory
  (ISIT)}, pages 511--515. IEEE, 2017.

\bibitem[Luo and Zhang(2020)]{luo2020tensor}
Yuetian Luo and Anru~R Zhang.
\newblock Tensor clustering with planted structures: Statistical optimality and
  computational limits.
\newblock \emph{arXiv preprint arXiv:2005.10743}, 2020.

\bibitem[Ma and Wu(2015)]{ma2015computational}
Zongming Ma and Yihong Wu.
\newblock Computational barriers in minimax submatrix detection.
\newblock \emph{The Annals of Statistics}, 43\penalty0 (3):\penalty0
  1089--1116, 2015.

\bibitem[Meka et~al.(2015)Meka, Potechin, and Wigderson]{meka2015sum}
Raghu Meka, Aaron Potechin, and Avi Wigderson.
\newblock Sum-of-squares lower bounds for planted clique.
\newblock In \emph{Proceedings of the forty-seventh annual ACM symposium on
  Theory of computing}, pages 87--96. ACM, 2015.

\bibitem[Richard and Montanari(2014)]{richard2014statistical}
Emile Richard and Andrea Montanari.
\newblock A statistical model for tensor {PCA}.
\newblock In \emph{Advances in Neural Information Processing Systems}, pages
  2897--2905, 2014.

\bibitem[Wang et~al.(2016)Wang, Berthet, and Samworth]{wang2016statistical}
Tengyao Wang, Quentin Berthet, and Richard~J Samworth.
\newblock Statistical and computational trade-offs in estimation of sparse
  principal components.
\newblock \emph{The Annals of Statistics}, 44\penalty0 (5):\penalty0
  1896--1930, 2016.

\bibitem[Wein et~al.(2019)Wein, El~Alaoui, and Moore]{wein2019kikuchi}
Alexander~S Wein, Ahmed El~Alaoui, and Cristopher Moore.
\newblock The kikuchi hierarchy and tensor pca.
\newblock In \emph{2019 IEEE 60th Annual Symposium on Foundations of Computer
  Science (FOCS)}, pages 1446--1468. IEEE, 2019.

\bibitem[Zhang and Xia(2018)]{zhang2018tensor}
Anru Zhang and Dong Xia.
\newblock Tensor {SVD}: Statistical and computational limits.
\newblock \emph{IEEE Transactions on Information Theory}, 64\penalty0
  (11):\penalty0 7311--7338, 2018.

\end{thebibliography}

\documentclass[final]{colt2020}
\usepackage{graphicx}
\usepackage{amssymb}
\usepackage{epstopdf}
\usepackage{natbib}
\usepackage{amssymb}
\usepackage{amsmath, mathabx}
\usepackage{mathrsfs}
\usepackage{hyperref}
\usepackage{comment}
\usepackage{xcolor}
\usepackage{enumitem}
\usepackage{algorithm}
\usepackage{algpseudocode}
\usepackage{stmaryrd}
\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}
\usepackage{fancyhdr}
\hypersetup{
    colorlinks=true,
    linkcolor=blue,
    filecolor=magenta,      
    urlcolor=cyan,
}
\newcommand{\A}{{\mathbf{A}}}

\newcommand{\bcA}{{\mathbfcal{A}}}

\newcommand{\bbP}{\mathbb{P}}

\newtheorem{Conjecture}{Conjecture}

\newtheorem{Challenge}{Challenge}

\DeclareMathAlphabet\mathbfcal{OMS}{cmsy}{b}{n}

\title{Open Problem: Average-Case Hardness of Hypergraphic Planted Clique Detection}%Is Hypergraphic Planted Clique Detection as hard as Planted Clique Detection?}
\coltauthor{%
 \Name{Yuetian Luo}$^\dagger$ \Email{yluo86@wisc.edu}
 \AND
 \Name{Anru R. Zhang}$^{\dagger *}$ \Email{anruzhang@stat.wisc.edu}\\
 \addr $^\dagger$Department of Statistics, University of Wisconsin-Madison\\ 
 $^*$Department of Biostatistics and Bioinformatics, Duke University
}

\date{\today}

\begin{document}
\maketitle

\setlength{\abovedisplayskip}{3pt}
\setlength{\belowdisplayskip}{3pt}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\begin{abstract}
We note the significance of hypergraphic planted clique (HPC) detection in the investigation of computational hardness for a range of tensor problems. We ask if more evidence for the computational hardness of HPC detection can be developed. In particular, we conjecture if it is possible to establish the equivalence of the computational hardness between HPC and PC detection. 
\end{abstract}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\section{Introduction}

The analysis of tensors has emerged as an active topic in recent years. Numerous datasets in the form of tensors emerge in various applications, such as collaborative filtering, neuroimaging analysis, hyperspectral imaging. In addition, tensor methods have been applied to many machine learning problems where the observations are not necessarily tensors, such as additive index models, high-order interaction pursuit, topic and latent variable models \citep{anandkumar2014tensor}. Compared with the inference for vectors or matrices, the tensor problems often possess distinct characteristics that bring significant computational challenges. As pointed out by the seminal work of \cite{hillar2013most}, 
extensions of many matrix concepts such as operator norm, singular values and eigenvalues are possible but computationally NP-hard. There often exist intrinsic gaps between the computational barriers and statistical limits in many tensor problems, as observed in tensor completion \citep{barak2016noisy}, tensor PCA \citep{richard2014statistical,hopkins2015tensor,wein2019kikuchi,zhang2018tensor,lesieur2017statistical,brennan2020reducibility}, high-order clustering \citep{luo2020tensor}, etc.  The analysis of computational barriers has attracted enormous attention because of its crucial role in the understanding of the computational feasibility of a wide range of tensor problems.

In the past decades, several lines of works have been devoted to the analysis of computational complexity in high-dimensional matrix problems. Such works include the average-case reduction \citep{berthet2013complexity}, analysis of restricted model of computation (e.g., statistical query \cite{kearns1998efficient}), analysis of algorithms such as sum-of-squares, belief propagation, approximate message passing, and analysis of optimization landscape. In particular, the average-case reduction requires randomized polynomial-time reduction from the conjectured hard problem in distribution to the target problem up to the conjectured computational barrier \citep{brennan2018reducibility}. Once the average-case reduction is established, all hardness results of the conjectured hard problem can be automatically inherited to the target problem. This provides a one-shot solution to the hardness of the target problem.

It would be ideal to use the same logic and establish the computational barriers for various tensor problems rigorously by average-case reduction from commonly raised conjectures, such as the planted clique (PC) detection or Boolean satisfiability problem (SAT), to the target tensor problem so that all of the hardness results of these well-studied conjectures could be naturally inherited. However, this route is often complicated by the multiway structure of the tensor. To overcome this, an alternative way of average-case reduction from hypergraphic planted clique (HPC) has been proposed in \cite{zhang2018tensor,luo2020tensor}. HPC has a more natural tensor structure that enables a more straightforward average-case reduction. A better understanding of the computational hardness relationship between PC detection and HPC detection can make a significant contribution to our knowledge of the average-case complexity in various tensor problems. 

\section{Planted Clique Detection and Hypergraphic Planted Clique Detection}
A $d$-hypergraph can be seen as an order-$d$ extension of regular graph. In a $d$-hypergraph $G = (V(G), E(G))$, each hyperedge $e \in E$ includes an unordered group of $d$ different vertices in $V$. Define $\mathcal{G}_d(N, 1/2)$ as the Erd\H{o}s-R{\'e}nyi $d$-hypergraph with $N$ vertices, where each hyperedge $(i_1, \ldots, i_d)$ is independently included in $E$ with probability $1/2$. Given a $d$-hypergraph $G$, define its adjacency tensor $\bcA = \bcA(G) \in (\{ 0,1 \}^N)^{\otimes d}$ as
\begin{equation*}
 	\bcA_{[i_1, \ldots, i_d]} = 
 	\begin{cases}
 	1, & \text{ if } (i_1, \ldots, i_d) \in E;\\
 	0, & \text{ otherwise}.	
 	\end{cases}
\end{equation*}
We define $\mathcal{G}_d(N, 1/2, \kappa)$ as the hypergraphic planted clique (HPC) model with the clique size $\kappa$. To generate $G \sim \mathcal{G}_d(N, 1/2, \kappa)$, we sample a random hypergraph from $\mathcal{G}_d(N, 1/2)$, pick $\kappa$ vertices uniformly at random from $[N]$, denote them as $K$, and connect all hyperedges $e$ if all vertices of $e$ are in $K$. The hypergraphic planted clique detection (HPC) can be formulated as the following hypothesis testing problem:
\begin{equation}\label{eq: HPC detection problem}
H_0: G \sim \mathcal{G}_d \left(N, 1/2\right)\quad \text{v.s.} \quad H_1: G \sim \mathcal{G}_d \left(N, 1/2, \kappa \right).
\end{equation}
We say a test $\phi$ solves the HPC detection if $\limsup_{N \to \infty} \bbP_{H_0} \left(\phi(\bcA) = 1\right) + \bbP_{H_1} \left(\phi(\bcA) = 0\right) = 0$, i.e., the sum of type-I, II errors converges to zero. 

When $d=2$, HPC detection reduces to the well-regard planted clique (PC) detection problem. The PC detection is statistically solvable if $\kappa \geq (2+\epsilon) \log_2 N$ for $\epsilon > 0$ by searching over all subgraphs of size $(2+\epsilon) \log_2 N$ \citep{bollobas1976cliques}. However, the best known polynomial-time algorithm requires $\kappa = \Omega(\sqrt{N})$ \citep{dekel2014finding,deshpande2015finding}. It has been widely conjectured that no polynomial-time algorithm can solve the PC detection when $\kappa = o(\sqrt{N})$. This conjecture has been strengthened in various types of computational models \citep{jerrum1992large,hopkins2018statistical,feldman2017statistical,deshpande2015improved,meka2015sum,barak2019nearly,gamarnik2019landscape}. As a result, the hardness conjecture of PC detection has become a key assumption in obtaining computational lower bounds for many problems, such as sparse PCA \citep{berthet2013complexity,wang2016statistical}, submatrix detection \citep{ma2015computational}, community detection \citep{hajek2015computational}, etc. 

Compared to PC detection, HPC detection ($d \geq 3$) has received far less attention in literature. \cite{bollobas1976cliques} proved that $K^d_N/\left(d!\log_2(N)\right)^{1/(d-1)} \overset{a.s.} \to 1$ if $K_N^d$ is the size of the largest fully connected subgraph in $G \sim \mathcal{G}_d(N, 1/2)$, which implies the HPC detection problem can be solved by exhaustive search in super-polynomial time when $\kappa \geq \left((d!+\epsilon) \log_2(N)\right)^{1/(d-1)}$ for any $\epsilon > 0$. \cite{zhang2018tensor} observed that applying SVD on the matricization/unfolding of $\bcA$ solves HPC detection efficiently if $\kappa = \Omega(\sqrt{N})$ but fails when $\kappa = N^{1/2 - \epsilon}$ for any $\epsilon>0$. Then they made a version of the following conjecture indicating that $\kappa = \Omega(\sqrt{N})$ is essential for solving HPC detection in polynomial time.
\begin{Conjecture}[HPC detection]\label{conj: hardness of tensor clique detection}
Suppose $d\geq 2$ is a fixed integer. If 
$$\limsup_{N \to \infty} \log \kappa / \log \sqrt{N} \leq 1 - \tau \quad \text{for any }\tau > 0,$$ 
for any sequence of polynomial-time tests $\{\phi \}_N: G \to \{0,1\}$, $\liminf_{N \to \infty} \bbP_{H_0} \left(\phi(G) = 1\right) + \bbP_{H_1} \left(\phi(G) = 0\right) > 1/2$. 
\end{Conjecture}
Recently, \cite{luo2020tensor} provides evidence for HPC detection conjecture by showing that \emph{Metropolis algorithms} \citep{jerrum1992large} and \emph{low-degree polynomial algorithms} \citep{hopkins2018statistical} can not solve the problem in polynomial-time when $\kappa = O(N^{\frac{1}{2} -\tau})$ for any $\tau > 0$.

\section{How hard is HPC detection?}

We raise the following challenges on the average-case computational hardness of HPC detection.

\vskip.1cm

\begin{Challenge}
	Can we establish the sharp limit of $\kappa$ to guarantee that HPC can be solved by SVD on the matricization/unfolding of $\bcA$? Is it possible to develop another polynomial-time algorithm for HPC that works below this limit?
\end{Challenge}

Such results for PC detection can be implied by \cite{knowles2013isotropic}  and \cite{deshpande2015finding}.

\begin{Challenge}
	 Can we provide more evidence for HPC detection conjecture under more classes of algorithms, such as statistical query (SQ), sum-of-squares (SOS), belief propagation (BP), approximate message passing (AMP), etc, in addition to the evidence in \cite{luo2020tensor}?
\end{Challenge}

\begin{Challenge}
	At the threshold level of $\kappa = O(N^{1/2-\tau})$ for $\tau > 0$, what is relationship between HPC detection and PC detection in their computational hardness? 
\end{Challenge}

First, HPC detection is no harder than PC detection. Given the adjacency tensor $\bcA$ of hypergraph $G$, by fixing $(d-2)$ indices and varying the remaining two, we get an adjacency matrix $\A$ from $\bcA$ which is a PC detection problem. So we can do a randomized polynomial-time reduction from HPC detection to PC detection without changing the planted clique size $\kappa$.

Now Challenge 3 becomes, how easier is HPC detection compared to PC detection? A one-shot approach to tackle this challenge is to do randomized polynomial-time reduction from PC to HPC. But this seems to be a very difficult task, upon which we have made several failed attempts. If reduction from PC to HPC could be proved, we would be able to arrive at the conclusion that PC and HPC detection are equivalently hard and all hardness results established in PC detection can be applied to HPC. The resulting connection between PC and HPC would enable solutions to a huge class of tensor computational hardness problems through existing matrix arguments. Such a connection can also provide solid evidence for the computational hardness assumptions of generalized planted clique problems proposed in \citep{brennan2020reducibility}, based on which we can establish various computational lower bounds more easily.

\bibliography{reference}

\end{document}